OpenStudy (anonymous):

Calculus: Verify that the function f(x) = 1/(x+1) satisfies the Mean Value Theorem on interval [0,2]. Then find all values of "c" such that f(2) - f(0) -------- 2 - 0 = f ' (c)

OpenStudy (anonymous):

\[ f'(x) = - \frac 1{(x+1)^2} \]

OpenStudy (anonymous):

My work so far: - f(x) is rational, and is continuous on {x|x =/= -1}. Then f(x) is continuous on [0,2]. - f' (x) = -1/(x+1)^2. f '(x) exists on {x|x =/= -1}. (MVT is fulfilled). - -1 / (c + 1)^2 = ( f(2) - f(0) ) / 2

OpenStudy (anonymous):

\[ \frac{1}{2} (f(2)-f(0))=-\frac{1}{3} = -\frac 1 {(1+c)^2 } \]

OpenStudy (anonymous):

wait. I think I know where I made my mistake I use the original function for f(2) and f(0)

OpenStudy (anonymous):

Solving the above gives you \[ c= -1 + \sqrt 3\\ c=-1-\sqrt 3 \] you take the positive root.